# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v2

@@ 2. The main result
The assumptions on coincidence of the associativity morphisms and braided symmetry as required in Theorem 8.19 are verified in the proof of Theorem 2.1 on the side of vertex operator algebras. For the associativity morphisms, methods are similar to those used to verify the assumption of Theorem 8.13. We apply Wenzl isomorphism and Drinfeld twist construction and then we use the structural properties of the Huang-Leposwky associativity morphisms in Theorem 31.9. For the special braiding morphisms we take into account the formulas given by Toledano-Laredo in Theorem 31.12. These two comparisons first motivated our method of proof for both braided symmetry and associativity morphisms.
The proof of (c) of Theorem 2.1 is reduced to the proof of (b) and application of Theorem 8.19. More in detail, for part (c) we use the fact that the special variables of the associativity morphisms and braided symmetry generate all the triples multiplicatively, in the sense that all the components except one in the variables of the associativity and braiding morphisms are taken by the generating representation in each Lie type. In the proof of Theorem 8.19, and the antecedent Theorems 8.13, we explain the generating property of these special values of the associativity and braiding morphisms with the pentagon equation of associativity morphisms and the hexagonal equations of a braided symmetry in a semisimple tensor category with a generating object but, as anticipated, another important property concerning the braid group representations in the category is needed, to determine the associativity morphisms uniquely.
11. The role of braid group duality for the generating object in determining the full braided tensor structure for the proof of (c).
The uniqueness Theorems 8.13 and 8.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory. The first instance of this property is Schur-Weyl duality for the general linear group. This duality says that the algebra of intertwining operators, from a tensor power of the vector representation $V$ of the general linear group to itself (also called centralizer algebras for the tensor powers of the vector representation) is generated by permutation operators [96].
@@
We need the property that the centralizer algebras $(V^{n},V^{n})$ with respect to the action of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ are generated by the natural representation of the braid group induced by the $R$-matrix. In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [91], [85], see also Definition 20.2 in [11]. The tensor powers $V^{n}$ appearing in our setting are canonical choices of truncations of the full tensor powers of the quantum group $U_{q}({\mathfrak{g}})$ obtained by induction over $n$.
The construction of the iterated truncations $V^{n}$ of $U_{q}({\mathfrak{g}})$ at the root of unity is explained in [91] and is the starting point to construct our weak Hopf algebras in Sections 29 and 30. The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 8.13 and 8.19. These theorems complete with cohomogical arguments what was proved up to that point in our work, coincidence of these structure maps on pairs and triples where one or two coordinated are fixed to the fundamental representation $V$. To do so, the way how our weak Hopf algebras are described in the application to the quantum group fusion category is also used.
@@
Returning to the question of whether the braid group representation generates the centralizer algebras $(V^{n},V^{n})$ for the truncated tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity, this question can be studied in the same way as for the case of a corresponding summand with $q$ not a root of unity, as in the truncated summand the root of unity property is not visible. This is explained among other things in Prop. 2.4 in [91], see also Prop 29.2 for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$. See also Section 3.5 in [91] for the special tensor products $V_{\lambda}\otimes V$ of representations of $U_{q}({\mathfrak{g}})$. Thus it suffices to have the duality property between the full algebras $(V^{\otimes n},V^{\otimes n})$ of $U_{q}({\mathfrak{g}})$ when $q$ is not a root of unity.
@@
The centralizer algebra of the fundamental representation for types $B$ and $D$ involve difficult work by Wenzl and his coauthors. Central theorems are Theorem 5.2 in [94] for the even case, and the previous Theorem 3.3 in [93] for the odd case. A special case has been studied by Rowell and Wenzl in [84].
@@
In the application of Theorems 8.13 and 8.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Section 30 are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
@@
Thus the fact that braid group duality for the fundamental representation $V$ was the relevant missing observation helpful to complete the proof of Theorem 2.1 in the previous version of [11], which led us to Theorems 8.13 and 8.19 as a completion of Lemma 7.10 in version 7 of [11], was for us a further important recognition of the relevance of the role of quantum groups. We refer the reader to the beautiful historical paper by V. F. R. Jones [57] for a general first introduction to the origin of quantum groups as solutions to the Yang-Baxter equation.
@@
We also would like to point out that our results of the categorical characterization of the class of unitary coboundary weak Hopf algebras that we obtain in Sect. 26, 27, 29 in [11] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [21], [22].
@@ 13. Outlook
 [^1] Claims, arguments, and reasoning of the recent papers [6], [7], [8] are essentially dependent on the results and in-depth discussions between S.C. and C.P. that took place from 2018 to 2022 on the contents of an entire project including [11] and the topics of the above papers, during a period in which S.C. was a co-author, before breaking the collaboration in early 2023 by his decision, with a minor contribution in a draft of a single section. In addition, C.P. department digital devices were subject to a serious hacking attack for several months in 2021 during the previous collaboration, which was recognized by authorities, that severely damaged her intellectual property and health, and the hacked material has not been restituted.
Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of affine vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem. She is also grateful to Marco V. Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [11] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper, and to Makoto Yamashita for interesting discussions on [40] and [92] during his visit at the math department of Rome, Sapienza, and interesting comments on our work. Finally she thanks Pavel Etingof for informing her of [29].

### v4

@@ 2. The main result
The assumptions on coincidence of the associativity morphisms and braided symmetry as required in Theorem 9.19 are verified in the proof of Theorem 2.1 on the side of vertex operator algebras. For the associativity morphisms, methods are similar to those used to verify the assumption of Theorem 9.13. We apply Wenzl isomorphism and Drinfeld twist construction and then we use the structural properties of the Huang-Leposwky associativity morphisms in Theorem 34.7 in [13]. For the special braiding morphisms we take into account the formulas given by Toledano-Laredo in Theorem 34.11 in [13]. These two comparisons first motivated our method of proof for both braided symmetry and associativity morphisms.
The proof of (d) of Theorem 2.1 is reduced to the proof of (c) and application of Theorem 9.19. More in detail, for part (d) we use the fact that the special variables of the associativity morphisms and braided symmetry generate all the triples multiplicatively, in the sense that all the components except one in the variables of the associativity and braiding morphisms are taken by the generating representation in each Lie type. In the proof of Theorem 9.19, and the antecedent Theorems 9.13, we explain the generating property of these special values of the associativity and braiding morphisms with the pentagon equation of associativity morphisms and the hexagonal equations of a braided symmetry in a semisimple tensor category with a generating object but, as anticipated, another important property concerning the braid group representations in the category is needed, to determine the associativity morphisms uniquely.
11. The role of braid group duality for the generating object in determining the full braided tensor structure for the proof of (d).
The uniqueness Theorems 9.13 and 9.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory. The first instance of this property is Schur-Weyl duality for the general linear group. This duality says that the algebra of intertwining operators, from a tensor power of the vector representation $V$ of the general linear group to itself (also called centralizer algebras for the tensor powers of the vector representation) is generated by permutation operators [104].
@@
We need the property that the centralizer algebras $(V^{n},V^{n})$ with respect to the action of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ are generated by the natural representation of the braid group induced by the $R$-matrix. In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [99], [92], see also Def. 20.1, 21.2 in [13]. The tensor powers $V^{n}$ appearing in our setting are canonical choices of truncations of the full tensor powers of the quantum group $U_{q}({\mathfrak{g}})$ obtained by induction over $n$.
The construction of the iterated truncations $V^{n}$ of $U_{q}({\mathfrak{g}})$ at the root of unity is explained in [99] and is the starting point to construct our weak Hopf algebras in Sect. 32. The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 9.13 and 9.19. These theorems complete with cohomogical arguments what was proved up to that point in our work, coincidence of these structure maps on pairs and triples where one or two coordinated are fixed to the fundamental representation $V$. To do so, the way how our weak Hopf algebras are described in the application to the quantum group fusion category is also used.
@@
Returning to the question of whether the braid group representation generates the centralizer algebras $(V^{n},V^{n})$ for the truncated tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity, this question can be studied in the same way as for the case of a corresponding summand with $q$ not a root of unity, as in the truncated summand the root of unity property is not visible. This is explained among other things in Prop. 2.4 in [99], see also Prop 31.2 in [13] for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$. See also Section 3.5 in [99] for the special tensor products $V_{\lambda}\otimes V$ of representations of $U_{q}({\mathfrak{g}})$. Thus it suffices to have the duality property between the full algebras $(V^{\otimes n},V^{\otimes n})$ of $U_{q}({\mathfrak{g}})$ when $q$ is not a root of unity.
@@
The centralizer algebra of the fundamental representation for types $B$ and $D$ involve difficult work by Wenzl and his coauthors. Central theorems are Theorem 5.2 in [102] for the even case, and the previous Theorem 3.3 in [101] for the odd case. A special case has been studied by Rowell and Wenzl in [91] in their study of a conjecture by Rowell and his coauthors regarding the characterization of finiteness of the image of the braid group in the morphism algebras of tensor powers of a simple object of a braided fusion category by integrality of the squared Frobenius-Perron dimension of the object.
@@
In the application of Theorems 9.13 and 9.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Sect. 32 in [13] are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
@@
Thus the fact that braid group duality for the fundamental representation $V$ was the relevant missing observation helpful to complete the proof of Theorem 2.1 in the previous version of [13], which led us to Theorems 9.13 and 9.19 as a completion of Lemma 7.10 in version 7 of [13], was for us a further important recognition of the relevance of the role of quantum groups. We refer the reader to the beautiful historical paper by V. F. R. Jones [64] for a general first introduction to the origin of quantum groups as solutions to the Yang-Baxter equation.
@@
We also would like to point out that our results of the categorical characterization of the class of unitary coboundary weak Hopf algebras that we obtain in Sect. 26, 27, 29 in [13] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [23], [24]. In Sect. 28 we explain the relation with the more familiar class of hermitian ribbon structures.
@@ 13. Outlook
 [^1] The author department digital devices were subject to a severe hacking attack for several months in 2021 which was recognized by authorities. The hacking severely damaged her intellectual property and health, and the hacked material has not been restituted. Related results have appeared in the papers [8], [9], [10]. These papers have been sued by the author to the Italian Procura della Repubblica between fall 2021 and spring 2023. Their claims, arguments, and reasoning essentially depend on [13], in-depth discussions between the author and S. Carpi that took place from 2018 to 2022 on the contents of an entire project including the topics of the above papers before his unilateral break of the collaboration in early 2023 and formal give up of his contribution, that consisted in a draft of a single section in 2018 based on [12], and yet unpublished joint work previously done by Sergio Ciamprone and the author (the first 60-page-long version of [13]), the original Finkelberg theorem, and work by Hans Wenzl in [99]. The author shared proofs of these attacks to several colleagues from 2022 to 2024.
Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem. She is also grateful to Marco V. Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [13] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper, and to Makoto Yamashita for interesting discussions on [42] and [100] during his visit at the math department of Rome, Sapienza, and interesting comments on our work. Finally she thanks Pavel Etingof for informing her of [31].